OpenStudy (anonymous):

What is the exact value of the expression sqrt(189) - sqrt(21) + sqrt(84) ? Simplify if possible.

OpenStudy (dumbsearch2):

I say it's 0. And that's an understatement.

OpenStudy (anonymous):

I have multiple choice, you want the answers?

OpenStudy (dumbsearch2):

I'll post them for you.

OpenStudy (dumbsearch2):

a √ 21 b 6 √ 7 c4√ 21 d24√ 7

OpenStudy (dumbsearch2):

The answer is sqrt(189) - sqrt(21) + sqrt(84) = sqrt(9 * 21) - sqrt(21) + sqrt(4 * 21) = 3 * sqrt(21) - sqrt(21) + 2 * sqrt(21) = 4 * sqrt(21)

OpenStudy (dumbsearch2):

Thanks to http://answers.yahoo.com/question/index?qid=20101226103625AAFvk0a Cheater. geez. Don't go there for the rest of the answers. It's just not right.

OpenStudy (dumbsearch2):

But give me a medal. Because you can. Click dat best answer button.

OpenStudy (anonymous):

Yay!! I guessed right